# Quality and Trust in LLM-generated Code

### 09 Feb 2024 v2

@@ Abstract.
Calibration has so far been studied in non-generative (e.g., classification) settings, especially in Software Engineering. However, generated code can quite often be wrong: Developers need to know when they should e.g., directly use, use after careful review, or discard model-generated code; thus Calibration is vital in generative settings. However, the notion of correctness of generated code is non-trivial, and thus so is Calibration. In this paper we make several contributions. We develop a framework for evaluating the Calibration of code-generating models. We consider several tasks, correctness criteria, datasets, and approaches, and find that by and large generative code models are not well-calibrated out of the box. We then show how Calibration can be improved, using standard methods such as Platt scaling. Our contributions will lead to better-calibrated decision-making in the current use of code generated by language models, and offers a framework for future research to further improve calibration methods for generative models in Software Engineering.
@@ 1. Introduction
Generative large language models (LLMs) are now widely-used for code completion in IDEs; in fact, LLMs have now reportedly become integrated into a wider range of software development activities including code review, code summarization, program repair, etc.. For example, Google’s Didact system (Large sequence models for software development activities, [n. d.]), supports debugging & repair, code review, and editing. However, LLMs come with documented risks: they might generate known buggy code (Jesse et al., 2023), or code with known & risky vulnerabilities (Asare et al., 2022; Schuster et al., 2021). The potential for bad code being generated by LLM “copilots” (Lo, 2023) is a widespread and growing risk to the software infrastructure that most modern societies rely on.
Nevertheless, current trends suggest that these LLM-based tools are becoming ever-more integrated into software development practices, potentially ushering in a future where most code is machine-written.
The motivation for our work is twofold: first, given that LLM-generated outputs could be wrong, humans who use such outputs would benefit from a clear and reliable indication of confidence that the generated code is correct. Second, such indications of confidence, when reliable and well-calibrated, can support sound engineering responses to possible quality issues, and thus improve the safety and reliability of the AI4SE ecosystem. Thus, for example, if an LLM offers a code suggestion, such a suggestion should ideally be accompanied by a numerical “confidence” signal, (e.g., a probability) indicating how likely the generated code is to match intent correctly. This signal should be well-calibrated: code suggestions with a high confidence score should usually be very good, and we should be able to reason about the likelihood and risks in a principled way. While calibration has been used in other settings (e.g. classically in weather prediction), we are the first to empirically study the calibration of generative large language models when used in software engineering settings, such as line-level code completion, function synthesis, and code-repair. When available, a well-calibrated measure would facilitate the development process, and eventually lead to higher code quality.
@@
  • We introduce the notion of calibration for confidence measures of LLM-generated code.
@@ 2.1.1. Calibration:
Calibration has deep roots in prediction problems like weather forecasting. Suppose a weather model predicts a 70% confidence (probability) of rain the next day. Is this confidence score helpful? As discussed in (Dawid, 1982), for a well-calibrated model, out of 100 days where the model predicted rain with 70% confidence, then in roughly 70 of those days it should have actually rained. That is, a model is well-calibrated if the model’s confidence in a given output, is quite close to the empirical relative frequency (likelihood) with which the output is actually correct. Good calibration supports further decision-making based on the confidence of a model. Poorly-calibrated, overconfident model—that often produces wrong output with high confidence—may lead to poor decisions; as might a diffident one that often outputs low-confidence when it is correct. A well-calibrated model informs the user exactly what the expectation of the prediction is. Thus, at varying degrees of confidence, a well-calibrated model gives a user the opportunity to respond in a calibrated way: at 20% confidence of rain, one might take a hat; at higher chances, one might carry an umbrella; at even higher chances, one might choose to take a bus rather than walk, etc. The idea of calibration has recently been a topic of interest in natural language processing. From an earlier work by (Jiang et al., 2021): given a model $M$, an input $X$ and true, expected correct output $Y$, a model output $M(X)=\hat{Y}$, (note that we won’t always have $Y=\hat{Y}$) and a output probability $P_{M}(\hat{Y}\,|\,X)$ provided by the model, a perfectly calibrated model satisfies the following condition (Adapted from (Jiang et al., 2021)):
@@
Usually, the level of calibration is imperfect; the yellow p above is typically not exactly the same as the first pink ${p}$. One empirically estimates the value of the yellow p from experimental samples, and measures the difference to the actual provided value. There are various measures of this difference, including Brier Score (Brier, 1950) and $ECE$ (Expected Calibration Error) (Pakdaman Naeini et al., 2015), which we describe below.
@@ 2.1.2. Why Calibration matters for Code
LLMs do make mistakes (Jesse et al., 2023), despite their wide use and popularity (Vaithilingam et al., 2022; Ziegler et al., 2022). To manage the risks of accepting bad code from an LLM, developers could use (Lo, 2023; Gros et al., 2023) a well-calibrated signal of confidence ${p}$ associated with generated code, for example. A well-calibrated high-value of ${p}$ would suggest a high empirical probabilty p that the code is correct. This sort of reliable signal of confidence in generated-code correctness supports better decision-making and proper use of code review resources, with well-placed trust and ready acceptance for code produced by the model with high-confidence of correctness, and appropriately careful review of code produced with less confidence. A poorly calibrated model may lead to either unnecessary review of likely correct code, or ill-advised acceptance of likely incorrect code. It should be noted here that good calibration allows more nuanced, effective quality-control (Q-C) decisions, beyond simple binary decisions (e.g. carefully review each token of generated code vs. just use it). For example, with a well-calibrated model, over 90% confidence might indicate that the generated code can be accepted with minimal review; below 10% confidence might indicate poor-quality code that should just be rejected; at 30% confidence the code might be subjected careful review, perhaps by more than one developer; at 70% perhaps the user themselves could review it, and accept it, and so on. This decision making also ties into reasoning about the risks of the domain. Such a well-calibrated quality-control process has been used in medicine e.g., for elder-care (Nierman et al., 2001), and for decision-making in cancer-care (Schwarz and Heider, 2019). While all these decision points should also consider the relative cost of these different Q-C decisions, a well-calibrated confidence signal is an important factor. Overall, given the cost & consequences of properly addressing software quality, and the potential benefits of LLM-generated code, a well-calibrated confidence signal is highly desirable.
@@ 3.1.1. Code Correctness
Unlike exact-match correctness test-passing correctness offers the advantage of admitting different semantically identical forms. However, test-passing correctness may be too permissive, or even wrong: test cases maybe insufficient or incorrect [^1] Recently, there has been a paper (Liu et al., 2023) accepted to NeurIPS, reporting inadequacies in the test sets of HumanEval; in addition tests can be “flaky” (Luo et al., 2014): the same test, on the same code, can sometimes pass and sometimes fail. In conclusion, both test-passing and exact-match afford automation, and are thus both appealing in an experimental setting such as ours; but both have limitations (see discussion of Table 4 below).
In this work, we studied both: we report test-passing correctness wherever tests are available, and when they are not, we report exact-match correctness. For the cases where tests were available, the exact-match results are presented in the appendix (see Table 7.
@@ 3.1.2. Confidence measures
(Lin et al., 2022; Zhou et al., 2023; Tian et al., 2023) (Reflective, $p_{v}$) We prompt the model to reflect jointly on the prompt, and the model-generated code, and then output a numerical value of its confidence in the generated code. An example is shown in Figure 1, on the left hand side. We use string-matching heuristics to handle percentages appearing different forms (e.g. “80%”, “80.00%”, “The probability is about 80.”, etc). For the line completion task, we experimented with a task-specific variation of the prompt, but found no improvement in rescaled calibration scores. Some extra logic is implemented for when the model fails to output a probability  [^3] Discussed further in appendix Understanding effects of verbalized retry failures.
@@
(Kadavath et al., 2022b) (Reflective, $p_{B}$ and $p_{NB}$) In this case, rather than prompting for a numerical score (as above), we prompt for a TRUE or FALSE answer. The probability associated with the TRUE token is taken to be the confidence measure. [^4] We transform raw Logits from the output layer of the LLM into probabilities using the softmax function. Additionally we extend this approach using normalization: the model can assign probability mass to multiple possible expressions of TRUE or FALSE or even other (e.g, “ True”, “ true”, “ ”, ). Thus when extending to the normalized form (Ask T/F N), we take the fraction of probability mass “True” assigned between only “True” and “False”. As GPT-3.5-Turbo-Instruct does not provide log-probabilities of the input tokens, we took the five most likely tokens for the first two completed tokens and selected the token with the maximum probability containing “true” or “false” (case insensitive), respectively.
@@
(Mixed) Other methods have been proposed (Zhang et al., 2023b). Self-consistency could also be used (Wang et al., 2022; Ahmed and Devanbu, 2023) could be another option. These are left for future research.
@@
For reflective measures, we expected that a code generation model should perform well (and provide well-calibrated confidence scores for correctness): first, for synthesis, given just a good natural language description (without tests), of the desired function; second, for completion & bug-fixing, given the surrounding context. In both settings, we measure correctness using available hidden test cases. We also note that including a failing test in an LLM prompt is not common experimental practice for the tasks we analyze. (Chen et al., 2021; Austin et al., 2021; Jiang et al., 2023).
@@ 3.1.3. Measures of Calibration
We use two measures of (mis)calibration: Brier Score (Brier, 1950) and $ECE$ (Pakdaman Naeini et al., 2015). Both measure the deviation from perfect calibration. As before (following (Guo et al., 2017)), we assume a model $M$, input $X$, actual desired output $Y$, and model prediction $M(X)=\hat{Y}$.
@@
where the model generates the code $M(x_{i})$ with confidence (probability) $\hat{p}_{i}$, and $\checkmark M(x_{i})$ indicates that the generated code is correct, as per the operative experimental definition. For a perfectly calibrated $M$, we would have $\text{corr}(S_{i})=\text{conf}(S_{i})\quad\forall i\in 1\ldots m$. In practice, we observe deviations from this ideal. Expected Calibration Error ($ECE$(Pakdaman Naeini et al., 2015) is a typical measure of calibration, calculated as the weighted average of these deviations.
@@
$ECE$ is intuitive, but can mislead; as seen below, a naive predictor whose confidence is always the base rate would yield a deceptively low $ECE$ value. An alternative measure of calibration, which avoids this issue is the Brier Score ${\mathcal{B}}$ (Brier, 1950); it
@@
Positive $SS$ (perfect score = 1.0) indicates improvement over baseline $\mathcal{B}_{ref}$; negative indicates making predictions worse than the baseline. Small positive values of $SS$ can sometimes indicate good skill; for example, the Deutsche Wetterdienst (German weather forecasting service) considers 0.05 Skill Score to be a minimum threshold for a good forecast quality  (Wetter und Klima, [n. d.]). As another data point, the American data journalism site 538 reports around a 0.13 skill in forecasting World Cup games (Jay Boice, [n. d.]), which is in the range of what we observe in our experiments for best case code generation by LLMs; but these LLM performance numbers are just a starting point, and can be expected to improve in the future. $ECE$ (Equation 6) and Brier Score (Equation 7) serve slightly different purposes: the Brier Score is calculated for each sample, and measures both the ability to correctly discriminate output categories, and calibration of the output probability. The $ECE$ measures just calibration; but it can be misleadingly low, as noted above for the unskilled predictor. Additionally binning must be done carefully, since it can affect $ECE$ scores (Nixon et al., 2019).
@@ 3.1.4. Rescaling Approaches
Machine learning models are not always calibrated. Guo et al (Guo et al., 2017) discuss ways of rescaling probability estimates to better match observations. A common approach is Platt scaling (Platt, 1999), where a logistic regression is fit to the logit values of the prediction i.e., the $\ln$ of the measured confidence probability. This optimizes two parameters, a linear scaling multiplier and a bias i.e., intercept, that shifts the value.
@@
Besides Platt scaling, temperature rescaling has also been used (Desai and Durrett, 2020; Kadavath et al., 2022a; Guo et al., 2017): this approach applies a scalar multiplier on the logits representing each class e.g., a multiclass image classifier. In our binary confidence case, this has similar expressivity to Platt scaling without an intercept. Other approaches include histogram binning (Zadrozny and Elkan, 2001), isotonic regression (Zadrozny and Elkan, 2002), inter alia (Guo et al., 2017; Kull et al., 2019). These approaches are more parameterized; given the data limitations in our experimental setup (e.g., a few hundred examples in function synthesis) pose higher risk of overfitting.
@@ 3.2.1. Function Synthesis
Figure 3a shows a sample prompt and model output. We use HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) datasets. One caveat: the samples in these datasets largely constitute artificial problems, specifically assembled to test the code-synthesis capacity of large language models; performance results (both accuracy and calibration) over these datasets may not necessarily generalize to real-world software development. Even so, these datasets provide a valuable datapoint for assessing model calibration.

### 16 Feb 2024 v3

# Calibration and Correctness of Language Models for Code

@@ Abstract.
Calibration has so far been studied in mostly non-generative (e.g., classification) settings, especially in Software Engineering. However, generated code can quite often be wrong: Developers need to know when they should e.g., directly use, use after careful review, or discard model-generated code; thus Calibration is vital in generative settings. However, the notion of correctness of generated code is non-trivial, and thus so is Calibration. In this paper we make several contributions. We develop a framework for evaluating the Calibration of code-generating models. We consider several tasks, correctness criteria, datasets, and approaches, and find that by and large generative code models are not well-calibrated out of the box. We then show how Calibration can be improved, using standard methods such as Platt scaling. Our contributions will lead to better-calibrated decision-making in the current use of code generated by language models, and offers a framework for future research to further improve calibration methods for generative models in Software Engineering.
@@ 1. Introduction
Generative large language models (LLMs) are now widely-used for code completion in IDEs; in fact, LLMs have now reportedly become integrated into a wider range of software development activities including code review, code summarization, program repair, etc.. For example, Google’s Didact system (researchLargeSequence, 35), supports debugging & repair, code review, and editing. However, LLMs come with documented risks: they might generate known buggy code (10174227, 25), or code with known & risky vulnerabilities (asare2022github, 3, 52). The potential for bad code being generated by LLM “copilots” (lo2023trustworthy, 40) is a widespread and growing risk to the software infrastructure that most modern societies rely on.
Nevertheless, current trends suggest that these LLM-based tools are becoming ever-more integrated into software development practices, potentially ushering in a future where most code is machine-written. The motivation for our work is twofold: first, given that LLM-generated outputs could be wrong, humans who use such outputs would benefit from a clear and reliable indication of confidence that the generated code is correct. Second, such indications of confidence, when reliable and well-calibrated, can support sound engineering responses to possible quality issues, and thus improve the safety and reliability of the AI4SE ecosystem. Thus, for example, if an LLM offers a code suggestion, such a suggestion should ideally be accompanied by a numerical “confidence” signal, (e.g., a probability) indicating how likely the generated code is to match intent correctly. This signal should be well-calibrated: code suggestions with a high confidence score should usually be very good, and we should be able to reason about the likelihood and risks in a principled way. While calibration has been used in other settings (e.g. classically in weather prediction), our focus here is to empirically study the calibration of generative large language models across practical software engineering settings, such as line-level code completion, function synthesis, and code-repair. When available, a well-calibrated measure would facilitate better risk-management in the development process, and eventually lead to higher code quality.
@@
  • We explore the notion of calibration for confidence measures of LLM-generated code.
@@ 2.1.1. Calibration:
Calibration has deep roots in prediction problems like weather forecasting. Suppose a weather model predicts a 70% confidence (probability) of rain the next day. Is this confidence score helpful? As discussed in (dawid1982well, 15), for a well-calibrated model, out of 100 days where the model predicted rain with 70% confidence, then in roughly 70 of those days it should have actually rained. That is, a model is well-calibrated if the model’s confidence in a given output, is quite close to the empirical relative frequency (likelihood) with which the output is actually correct. Good calibration supports further decision-making based on the confidence of a model. Poorly-calibrated, overconfident model—that often produces wrong output with high confidence—may lead to poor decisions; as might a diffident one that often outputs low-confidence when it is correct. A well-calibrated model informs the user exactly what the expectation of the prediction is. Thus, at varying degrees of confidence, a well-calibrated model gives a user the opportunity to respond in a calibrated way: at 20% confidence of rain, one might take a hat; at higher chances, one might carry an umbrella; at even higher chances, one might choose to take a bus rather than walk, etc. The idea of calibration has recently been a topic of interest in natural language processing. From an earlier work by (jiang2021can, 27): given a model $M$, an input $X$ and true, expected correct output $Y$, a model output $M(X)=\hat{Y}$, (note that we won’t always have $Y=\hat{Y}$) and a output probability $P_{M}(\hat{Y}\,|\,X)$ provided by the model, a perfectly calibrated model satisfies the following condition (Adapted from (jiang2021can, 27)):
@@
Usually, the level of calibration is imperfect; the yellow p above is typically not exactly the same as the first pink ${p}$. One empirically estimates the value of the yellow p from experimental samples, and measures the difference to the actual provided value. There are various measures of this difference, including Brier Score (brier1950verification, 9) and $ECE$ (Expected Calibration Error) (Pakdaman_Naeini_Cooper_Hauskrecht_2015, 48), which we describe below.
@@ 2.1.2. Why Calibration matters for Code
LLMs do make mistakes (10174227, 25), despite their wide use and popularity (vaithilingam2022expectation, 59, 71). To manage the risks of accepting bad code from an LLM, developers could use (lo2023trustworthy, 40, 19) a well-calibrated signal of confidence ${p}$ associated with generated code, for example. A well-calibrated high-value of ${p}$ would suggest a high empirical probabilty p that the code is correct. This sort of reliable signal of confidence in generated-code correctness supports better decision-making and proper use of code review resources, with well-placed trust and ready acceptance for code produced by the model with high-confidence of correctness, and appropriately careful review of code produced with less confidence. A poorly calibrated model may lead to either unnecessary review of likely correct code, or ill-advised acceptance of likely incorrect code. It should be noted here that good calibration allows more nuanced, effective quality-control (Q-C) decisions, beyond simple binary decisions (e.g. carefully review each token of generated code vs. just use it). For example, with a well-calibrated model, over 90% confidence might indicate that the generated code can be accepted with minimal review; below 10% confidence might indicate poor-quality code that should just be rejected; at 30% confidence the code might be subjected careful review, perhaps by more than one developer; at 70% perhaps the user themselves could review it, and accept it, and so on. This decision making also ties into reasoning about the risks of the domain. Such a well-calibrated quality-control process has been used in medicine e.g., for elder-care (nierman2001outcome, 44), and for decision-making in cancer-care (schwarz2019guess, 53). While all these decision points should also consider the relative cost of these different Q-C decisions, a well-calibrated confidence signal is an important factor. Overall, given the cost & consequences of properly addressing software quality, and the potential benefits of LLM-generated code, a well-calibrated confidence signal is highly desirable.
@@ 3.1.1. Code Correctness
Unlike exact-match correctness test-passing correctness offers the advantage of admitting different semantically identical forms. However, test-passing correctness may be too permissive, or even wrong: test cases maybe insufficient or incorrect [^1] Recently, there has been a paper (liu2023code, 39) accepted to NeurIPS, reporting inadequacies in the test sets of HumanEval; in addition tests can be “flaky” (luo2014empirical, 42): the same test, on the same code, can sometimes pass and sometimes fail. In conclusion, both test-passing and exact-match afford automation, and are thus both appealing in an experimental setting such as ours; but both have limitations (see discussion of Table 4 below).
In this work, we studied both: we report test-passing correctness wherever tests are available, and when they are not, we report exact-match correctness. For the cases where tests were available, the exact-match results are presented in the appendix (see Table 7).
@@ 3.1.2. Confidence measures
(lin2022teaching, 38, 70, 58) (Reflective, $p_{v}$) We prompt the model to reflect jointly on the prompt, and the model-generated code, and then output a numerical value of its confidence in the generated code. An example is shown in Figure 1, on the left hand side. We use string-matching heuristics to handle percentages appearing different forms (e.g. “80%”, “80.00%”, “The probability is about 80.”, etc). For the line completion task, we experimented with a task-specific variation of the prompt, but found no improvement in rescaled calibration scores. Some extra logic is implemented for when the model fails to output a probability  [^3] Discussed further in appendix Understanding effects of verbalized retry failures.
@@
(kadavath2022language, 30) (Reflective, $p_{B}$ and $p_{NB}$) In this case, rather than prompting for a numerical score (as above), we prompt for a TRUE or FALSE answer. The probability associated with the TRUE token is taken to be the confidence measure. [^4] We transform raw Logits from the output layer of the LLM into probabilities using the softmax function. Additionally we extend this approach using normalization: the model can assign probability mass to multiple possible expressions of TRUE or FALSE or even other (e.g, “ True”, “ true”, “ ”, ). Thus when extending to the normalized form (Ask T/F N), we take the fraction of probability mass “True” assigned between only “True” and “False”. As GPT-3.5-Turbo-Instruct does not provide log-probabilities of the input tokens, we took the five most likely tokens for the first two completed tokens and selected the token with the maximum probability containing “true” or “false” (case insensitive), respectively.
@@
(Mixed) Other methods have been proposed (zhang2023coder, 68, 32). Self-consistency could also be used (wang2022self, 60, 1) could be another option. These are left for future research.
@@
For reflective measures, we expected that a code generation model should perform well (and provide well-calibrated confidence scores for correctness): first, for synthesis, given just a good natural language description (without tests), of the desired function; second, for completion & bug-fixing, given the surrounding context. In both settings, we measure correctness using available hidden test cases. We also note that including a failing test in an LLM prompt is not common experimental practice for the tasks we analyze. (chen2021evaluating, 12, 4, 26).
@@ 3.1.3. Measures of Calibration
We use two measures of (mis)calibration: Brier Score (brier1950verification, 9) and $ECE$ (Pakdaman_Naeini_Cooper_Hauskrecht_2015, 48). Both measure the deviation from perfect calibration. As before (following (guo2017calibration, 20)), we assume a model $M$, input $X$, actual desired output $Y$, and model prediction $M(X)=\hat{Y}$.
@@
where the model generates the code $M(x_{i})$ with confidence (probability) $\hat{p}_{i}$, and $\checkmark M(x_{i})$ indicates that the generated code is correct, as per the operative experimental definition. For a perfectly calibrated $M$, we would have $\text{corr}(S_{i})=\text{conf}(S_{i})\quad\forall i\in 1\ldots m$. In practice, we observe deviations from this ideal. Expected Calibration Error ($ECE$(Pakdaman_Naeini_Cooper_Hauskrecht_2015, 48) is a typical measure of calibration, calculated as the weighted average of these deviations.
@@
$ECE$ is intuitive, but can mislead; as seen below, a naive predictor whose confidence is always the base rate would yield a deceptively low $ECE$ value. An alternative measure of calibration, which avoids this issue is the Brier Score ${\mathcal{B}}$ (brier1950verification, 9); it
@@
Positive $SS$ (perfect score = 1.0) indicates improvement over baseline $\mathcal{B}_{ref}$; negative indicates making predictions worse than the baseline. Small positive values of $SS$ can sometimes indicate good skill; for example, the Deutsche Wetterdienst (German weather forecasting service) considers 0.05 Skill Score to be a minimum threshold for a good forecast quality  (dwdWetterKlima, 62). As another data point, the American data journalism site 538 reports around a 0.13 skill in forecasting World Cup games (fivethirtyeightGoodFiveThirtyEight, 24), which is in the range of what we observe in our experiments for best case code generation by LLMs; but these LLM performance numbers are just a starting point, and can be expected to improve in the future. $ECE$ (Equation 6) and Brier Score (Equation 7) serve slightly different purposes: the Brier Score is calculated for each sample, and measures both the ability to correctly discriminate output categories, and calibration of the output probability. The $ECE$ measures just calibration; but it can be misleadingly low, as noted above for the unskilled predictor. Additionally binning must be done carefully, since it can affect $ECE$ scores (nixon2019measuring, 47).
@@ 3.1.4. Rescaling Approaches
Machine learning models are not always calibrated. Guo et al (guo2017calibration, 20) discuss ways of rescaling probability estimates to better match observations. A common approach is Platt scaling (Platt1999ProbabilisticOF, 50), where a logistic regression is fit to the logit values of the prediction i.e., the $\ln$ of the measured confidence probability. This optimizes two parameters, a linear scaling multiplier and a bias i.e., intercept, that shifts the value.
@@
Besides Platt scaling, temperature rescaling has also been used (desai2020calibration, 16, 29, 20): this approach applies a scalar multiplier on the logits representing each class e.g., a multiclass image classifier. In our binary confidence case, this has similar expressivity to Platt scaling without an intercept. Other approaches include histogram binning (Zadrozny2001ObtainingCP, 63), isotonic regression (Zadrozny2002TransformingCS, 64), inter alia (guo2017calibration, 20, 34). These approaches are more parameterized; given the data limitations in our experimental setup (e.g., a few hundred examples in function synthesis) pose higher risk of overfitting.
@@ 3.2.1. Function Synthesis
Figure 3a shows a sample prompt and model output. We use HumanEval (chen2021evaluating, 12) and MBPP (austin2021program, 4) datasets. One caveat: the samples in these datasets largely constitute artificial problems, specifically assembled to test the code-synthesis capacity of large language models; performance results (both accuracy and calibration) over these datasets may not necessarily generalize to real-world software development. Even so, these datasets provide a valuable datapoint for assessing model calibration.